# The Hot Jupiter Radius Anomaly and Stellar Connections

### v1

@@ 6.3 Magnetic Fields
The s of exoplanets are poorly understood due to the difficulty in making direct measurements of them, but is thought to relate to the rotation rate, intrinsic temperature, or both (Reiners and Christensen 2010). Hot Jupiters are expected to be tidally locked Jackson et al. (2008), so even on their short periods the rotation rates will be lower than those of Jupiter and Saturn. On the other hand, their intrinsic temperatures appear to be substantially higher (Sec. 6.2). Assuming that hot Jupiter radii are explained by injected heat, Yadav and Thorngren (2017) applied the Christensen et al. (2009) scaling relations to infer hot Jupiter magnetic field strengths. The mean field on the dynamo surface from Christensen et al. (2009) was:
@@
An important angle for observationally constraining hot Jupiter magnetic fields is the interactions between the fields of the planet and its parent star. The is known to be a tracer of magnetic activity in stellar chromospheres (e.g., Schrijver et al. 1989), and so it has been hypothesized as a potential avenue for measuring star-planet magnetic interactions (see Lanza 2008; Poppenhaeger 2019), though the point of peak activity may not be directly facing the planet. Subsequent observations have found cases where the Ca II K line varies at the orbital period of the planet (e.g., Shkolnik et al. 2003, 2005), supporting the hypothesis that it is driven by star-planet interactions. An important development was made by Cauley et al. (2019), who calculated the magnetic field strengths implied by the Ca II K line modulation amplitudes for four hot Jupiters and found strong magnetic fields with surface strengths of 20-120 G, consistent (if possibly slightly lower) with the intrinsic temperature-based scaling laws used by Yadav and Thorngren (2017). Thus it provides observational support for the hypothesis that hot Jupiter magnetic fields are significantly stronger than those of cooler giants like Jupiter and Saturn.
@@ 6.4 Tidal Effects
A case of particular interest is that of HAT-P-13 b and c. Using a secular dynamics analysis, (Batygin et al. 2009) showed that the system fairly quickly reaches a steady-state in which gravitational interactions between b and c transfer orbital energy into the eccentricity of b, which is then damped by circularization tides onto a constant value of . More importantly, the two orbits precess together, and while the details of the secular dynamics arguments are too involved to go into here, the result is that the eccentricity of b is directly related to it’s value of ! Follow-up by (Buhler et al. 2016) used HST secondary transits to determine an eccentricity of , implying , which corresponds to a most likely core mass of and a 1-sigma upper limit of , reasonable numbers for the object. This is so far the only constraint on an exoplanet’s precise enough to be useful for constraining its interior structure, but it seems likely that more will be measured, especially as more long-period companions of known exoplanets are discovered.
@@ 7 Conclusion
The hot Jupiter radius inflation problem continues to be one of the longest-standing open questions in exoplanet science; however, over the last two decades we have made significant progress in collecting observational evidence for its cause. To review the key points of evidence from previous sections:
@@
The absence of cold traps on WASP-121 b indicates a high intrinsic temperature (Sing et al. 2019b).
@@
Alternatively, we may need to consider multiple mechanisms operating at the same time. If both thermal tides and Ohmic dissipation contribute significantly, for example, they might match the observations by having Ohmic dissipation regulate heat flow out of the planet while modest thermal tides allow the planet to reinflate. The key point here is that we shouldn’t require that any one inflation mechanism solve the entire problem on its own – they may be acting in concert with one another.
Let’s conclude with some promising areas of investigation. On the observational side, the discovery of more hot Jupiters is always helpful for doing population statistics, but those at very young ages or around post-main sequence stars will be particularly valuable as constraints on the radius evolution of hot Jupiters. For individual planet studies, any atmosphere observations that can measure or constrain the intrinsic temperature (as in Sing et al. 2019b) are of particular interest, as would constraints on the atmospheric metallicity, as these help to narrow down the range of plausible interiors. Observations of the magnetic field strengths of hot Jupiters is similarly helpful, be they through star-planet interactions, direct radio observation, or some other method. Finally, a sufficiently precise observation of the prolate shape of a hot Jupiter (or equivalently its Love number) would provide a valuable new angle of constraint on the deep interior.
@@
Between the abundance of new observations and the ongoing theoretical work to understand them, this is a very exciting time for the study of hot Jupiters. The coming years will likely see further dramatic improvements in our understanding of these planets.
@@ Acknowledgements.
- Adams (2011) Adams FC (2011) Magnetically Controlled Outflows from Hot Jupiters. The Astrophysical Journal 730(1):27
@@
- Bastian et al. (2018) Bastian TS, Villadsen J, Maps A, Hallinan G Beasley AJ (2018) Radio Emission from the Exoplanetary System Eridani. The Astrophysical Journal 857(2):133
- Batygin and Stevenson (2010) Batygin K Stevenson DJ (2010) Inflating Hot Jupiters with Ohmic Dissipation. The Astrophysical Journal Letters 714(2):L238
@@
- Beltz et al. (2023) Beltz H, Rauscher E, Kempton EMR, Malsky I Savel AB (2023) Magnetic Effects and 3D Structure in Theoretical High-resolution Transmission Spectra of Ultrahot Jupiters: The Case of WASP-76b. The Astronomical Journal 165(6):257
@@
- Burrows et al. (2008) Burrows A, Ibgui L Hubeny I (2008) Optical Albedo Theory of Strongly Irradiated Giant Planets: The Case of HD 209458b. The Astrophysical Journal 682(2):1277–1282
@@
- Chabrier et al. (2004) Chabrier G, Barman T, Baraffe I, Allard F Hauschildt PH (2004) The Evolution of Irradiated Planets: Application to Transits. The Astrophysical Journal 603(1):L53–L56
@@
- Demory and Seager (2011) Demory BO Seager S (2011) Lack of Inflated Radii for Kepler Giant Planet Candidates Receiving Modest Stellar Irradiation. The Astrophysical Journal Supplement Series 197(1):12
- Dos Santos (2021) Dos Santos LA (2021) Observations of planetary winds and outflows. Proceedings of the International Astronomical Union 17(S370):56–71
@@
- Fortney et al. (2010) Fortney JJ, Shabram M, Showman AP et al. (2010) Transmission Spectra of Three-Dimensional Hot Jupiter Model Atmospheres. The Astrophysical Journal 709:1396–1406
@@
- Fuentes and Cumming (2020) Fuentes JR Cumming A (2020) Penetration of a cooling convective layer into a stably-stratified composition gradient: Entrainment at low Prandtl number. arXiv:200704265 [astro-ph, physics:physics]
@@
- Haldemann et al. (2020) Haldemann J, Alibert Y, Mordasini C Benz W (2020) AQUA: A Collection of H$_2$O Equations of State for Planetary Models. Astronomy & Astrophysics 643:A105
@@
- Huang and Cumming (2012) Huang X Cumming A (2012) Ohmic Dissipation in the Interiors of Hot Jupiters. The Astrophysical Journal 757(1):47
@@
- Jackson et al. (2008) Jackson B, Greenberg R Barnes R (2008) Tidal Heating of Extrasolar Planets. The Astrophysical Journal 681:1631–1638
@@
- Kao et al. (2018) Kao MM, Hallinan G, Pineda JS, Stevenson D Burgasser A (2018) The Strongest Magnetic Fields on the Coolest Brown Dwarfs. The Astrophysical Journal Supplement Series 237(2):25
@@
- Lanza (2008) Lanza AF (2008) Hot Jupiters and stellar magnetic activity. Astronomy & Astrophysics 487(3):1163–1170
@@
- Leconte and Chabrier (2012) Leconte J Chabrier G (2012) A new vision of giant planet interiors: Impact of double diffusive convection. Astronomy & Astrophysics 540:A20
- Leconte et al. (2010) Leconte J, Chabrier G, Baraffe I Levrard B (2010) Is tidal heating sufficient to explain bloated exoplanets? Consistent calculations accounting for finite initial eccentricity. Astronomy & Astrophysics 516:A64
- Leconte et al. (2011) Leconte J, Chabrier G, Baraffe I Levrard B (2011) The radius anomaly in the planet/brown dwarf overlapping mass regime. In: European Physical Journal Web of Conferences, eprint: arXiv:1011.0336, vol 11, p 03004, DOI 10.1051/epjconf/20101103004
@@
- Lee et al. (2022) Lee EJ, Fuentes JR Hopkins PF (2022) Establishing Dust Rings and Forming Planets within Them. The Astrophysical Journal 937:95
@@
- Love (1909) Love AEH (1909) The yielding of the earth to disturbing forces. Proceedings of the Royal Society of London Series A, Containing Papers of a Mathematical and Physical Character 82(551):73–88
@@
- Mankovich and Fuller (2021) Mankovich C Fuller J (2021) A diffuse core in Saturn revealed by ring seismology. arXiv:210413385 [astro-ph]
@@
- Marley et al. (2007) Marley MS, Fortney JJ, Hubickyj O, Bodenheimer P Lissauer JJ (2007) On the Luminosity of Young Jupiters. The Astrophysical Journal 655(1):541
@@
- May et al. (2021) May EM, Komacek TD, Stevenson KB et al. (2021) Spitzer Phase-curve Observations and Circulation Models of the Inflated Ultrahot Jupiter WASP-76b. The Astronomical Journal 162(4):158
@@
- Millholland and Spalding (2020) Millholland SC Spalding C (2020) Formation of Ultra-short-period Planets by Obliquity-driven Tidal Runaway. The Astrophysical Journal 905:71
@@
- Naef et al. (2001) Naef D, Latham DW, Mayor M et al. (2001) HD 80606 b, a planet on an extremely elongated orbit. Astronomy & Astrophysics 375(2):L27–L30
- Narang et al. (2023) Narang M, Oza AV, Hakim K et al. (2023) Radio-loud Exoplanet-exomoon Survey: GMRT Search for Electron Cyclotron Maser Emission. The Astronomical Journal 165(1):1
@@
- Parmentier et al. (2016) Parmentier V, Fortney JJ, Showman AP, Morley C Marley MS (2016) Transitions in the Cloud Composition of Hot Jupiters. The Astrophysical Journal 828(1):22
- Paxton et al. (2013) Paxton B, Cantiello M, Arras P et al. (2013) Modules for Experiments in Stellar Astrophysics (MESA): Planets, Oscillations, Rotation, and Massive Stars. The Astrophysical Journal Supplement Series 208(1):4
- Piso et al. (2015) Piso AMA, Youdin AN Murray-Clay RA (2015) Minimum Core Masses for Giant Planet Formation With Realistic Equations of State and Opacities. The Astrophysical Journal 800(2):82
@@
- Poppenhaeger (2019) Poppenhaeger K (2019) How stars and planets interact: A look through the high-energy window. Astronomische Nachrichten 340(4):329–333
@@
- Reiners and Christensen (2010) Reiners A Christensen UR (2010) A magnetic field evolution scenario for brown dwarfs and giant planets. Astronomy & Astrophysics 522:A13
- Ricker et al. (2014) Ricker GR, Winn JN, Vanderspek R et al. (2014) Transiting Exoplanet Survey Satellite. Journal of Astronomical Telescopes, Instruments, and Systems 1(1):014,003
@@
- Sainsbury-Martinez et al. (2019) Sainsbury-Martinez F, Wang P, Fromang S et al. (2019) Idealised simulations of the deep atmosphere of hot Jupiters: Deep, hot adiabats as a robust solution to the radius inflation problem. Astronomy & Astrophysics 632:A114
@@
- Schrijver et al. (1989) Schrijver CJ, Cote J, Zwaan C Saar SH (1989) Relations between the photospheric magnetic field and the emission from the outer atmospheres of cool stars. I - The solar CA II K line core emission. The Astrophysical Journal 337:964
- Serenelli and Basu (2010) Serenelli AM Basu S (2010) Determining the Initial Helium Abundance of the Sun. The Astrophysical Journal 719(1):865
@@
- Shkolnik et al. (2003) Shkolnik E, Walker GAH Bohlender DA (2003) Evidence for Planet-induced Chromospheric Activity on HD 179949. The Astrophysical Journal 597(2):1092–1096
- Shkolnik et al. (2005) Shkolnik E, Walker GAH, Bohlender DA, Gu PG Kurster M (2005) Hot Jupiters and Hot Spots: The Short- and Long-Term Chromospheric Activity on Stars with Giant Planets. The Astrophysical Journal 622(2):1075–1090
- Showman and Polvani (2011) Showman AP Polvani LM (2011) Equatorial Superrotation on Tidally Locked Exoplanets. The Astrophysical Journal 738(1):71
- Sing et al. (2019a) Sing DK, Lavvas P, Ballester GE et al. (2019a) The Hubble Space Telescope PanCET Program: Exospheric Mg II and Fe II in the Near-ultraviolet Transmission Spectrum of WASP-121b Using Jitter Decorrelation. The Astronomical Journal 158:91
- Sing et al. (2019b) Sing DK, Lavvas P, Ballester GE et al. (2019b) The Hubble Space Telescope PanCET Program: Exospheric Mg ii and Fe ii in the Near-ultraviolet Transmission Spectrum of WASP-121b Using Jitter Decorrelation. The Astronomical Journal 158(2):91
@@
- Thorngren et al. (2019b) Thorngren DP, Marley MS Fortney JJ (2019b) An Empirical Mass-Radius Relation for Cool Giant Planets. arXiv e-prints 1909:arXiv:1909.09,207
@@
- Valsecchi et al. (2014) Valsecchi F, Rasio FA Steffen JH (2014) From Hot Jupiters to Super-Earths via Roche Lobe Overflow. The Astrophysical Journal 793(1):L3
@@
- Vazan et al. (2018) Vazan A, Ormel CW Dominik C (2018) Effect of core cooling on the radius of sub-Neptune planets. Astronomy and Astrophysics 610:L1
@@
- Wong et al. (2021) Wong I, Shporer A, Zhou G et al. (2021) TOI-2109: An Ultrahot Gas Giant on a 16 hr Orbit. The Astronomical Journal 162(6):256
@@
- Yee et al. (2021) Yee SW, Winn JN Hartman JD (2021) How Complete Are Surveys for Nearby Transiting Hot Jupiters? The Astronomical Journal 162(6):240
@@
- Zarka (1998) Zarka P (1998) Auroral radio emissions at the outer planets: Observations and theories. Journal of Geophysical Research: Planets 103(E9):20,159–20,194

### v3

@@ 6.3 Magnetic Fields
The magnetic fields of exoplanets are poorly understood due to the difficulty in making direct measurements of them, but are thought to relate to the rotation rate, intrinsic temperature, or both (Reiners and Christensen 2010). Hot Jupiters are expected to be tidally locked Jackson et al. (2008), so even on their short periods the rotation rates will be lower than those of Jupiter and Saturn. On the other hand, their intrinsic temperatures appear to be substantially higher (Sec. 6.2). Assuming that hot Jupiter radii are explained by injected heat, Yadav and Thorngren (2017) applied the Christensen et al. (2009) scaling relations to infer hot Jupiter magnetic field strengths. The mean field on the dynamo surface from Christensen et al. (2009) was:
@@
An important angle for observationally constraining hot Jupiter magnetic fields is the interactions between the fields of the planet and its parent star. The Ca II K line is known to be a tracer of magnetic activity in stellar chromospheres (e.g., Schrijver et al. 1989), and so it has been hypothesized as a potential avenue for measuring star-planet magnetic interactions (see Lanza 2008; Poppenhaeger 2019), though the point of peak activity may not be directly facing the planet. Subsequent observations have found cases where the Ca II K line varies at the orbital period of the planet (e.g., Shkolnik et al. 2003, 2005), supporting the hypothesis that it is driven by star-planet interactions. An important development was made by Cauley et al. (2019), who calculated the magnetic field strengths implied by the Ca II K line modulation amplitudes for four hot Jupiters and found strong magnetic fields with surface strengths of 20-120 G, consistent (if possibly slightly lower) with the intrinsic temperature-based scaling laws used by Yadav and Thorngren (2017). Thus it provides observational support for the hypothesis that hot Jupiter magnetic fields are significantly stronger than those of cooler giants like Jupiter and Saturn.
@@ 6.4 Tidal Effects
A case of particular interest is that of HAT-P-13 b and c. Using a secular dynamics analysis, (Batygin et al. 2009) showed that the system fairly quickly reaches a steady-state in which gravitational interactions between b and c transfer orbital energy into the eccentricity of b, which is then damped by circularization tides onto a constant value of . More importantly, the two orbits precess together, and while the details of the secular dynamics arguments are too involved to go into here, the result is that the eccentricity of b is directly related to its value of ! Follow-up by (Buhler et al. 2016) used HST secondary transits to determine an eccentricity of , implying , which corresponds to a most likely core mass of and a 1-sigma upper limit of , reasonable numbers for the object. This is so far the only constraint on an exoplanet’s precise enough to be useful for constraining its interior structure, but it seems likely that more will be measured, especially as more long-period companions of known exoplanets are discovered.
@@ 7 Conclusion
The hot Jupiter radius inflation problem continues to be one of the longest-standing open questions in exoplanet science; however, significant progress has been made over the last two decades in collecting observational evidence for its cause. To review the key points of evidence from previous sections:
@@
The absence of cold traps on WASP-121 b indicates a high intrinsic temperature (Sing et al. 2019).
@@
Alternatively, it may be necessary to consider multiple mechanisms operating at the same time. If both thermal tides and Ohmic dissipation contribute significantly, for example, they might match the observations by having Ohmic dissipation regulate heat flow out of the planet while modest thermal tides allow the planet to reinflate. The key point here is that it isn’t necessarily the case that any one inflation mechanism solve the entire problem on its own – they may be acting in concert with one another.
To conclude, here are some promising areas of investigation. On the observational side, the discovery of more hot Jupiters is always helpful for doing population statistics, but those at very young ages or around post-main sequence stars will be particularly valuable as constraints on the radius evolution of hot Jupiters. For individual planet studies, any atmosphere observations that can measure or constrain the intrinsic temperature (as in Sing et al. 2019) are of particular interest, as would constraints on the atmospheric metallicity, as these help to narrow down the range of plausible interiors. Observations of the magnetic field strengths of hot Jupiters is similarly helpful, be they through star-planet interactions, direct radio observation, or some other method. Finally, a sufficiently precise observation of the prolate shape of a hot Jupiter (or equivalently its Love number) would provide a valuable new angle of constraint on the deep interior.
@@
Between the abundance of new observations and the ongoing theoretical work to understand them, this is a very exciting time for the study of hot Jupiters. The coming years will likely see further dramatic advancements in understanding of these planets.
@@ Acknowledgements.
- Adams (2011) Adams FC (2011) Magnetically Controlled Outflows from Hot Jupiters. ApJ 730(1):27
@@
- Bastian et al. (2018) Bastian TS, Villadsen J, Maps A, Hallinan G Beasley AJ (2018) Radio Emission from the Exoplanetary System Eridani. ApJ 857(2):133
- Batygin and Stevenson (2010) Batygin K Stevenson DJ (2010) Inflating Hot Jupiters with Ohmic Dissipation. ApJL 714(2):L238
@@
- Beltz et al. (2023) Beltz H, Rauscher E, Kempton EMR, Malsky I Savel AB (2023) Magnetic Effects and 3D Structure in Theoretical High-resolution Transmission Spectra of Ultrahot Jupiters: The Case of WASP-76b. AJ 165(6):257
@@
- Burrows et al. (2008) Burrows A, Ibgui L Hubeny I (2008) Optical Albedo Theory of Strongly Irradiated Giant Planets: The Case of HD 209458b. ApJ 682(2):1277–1282
@@
- Chabrier et al. (2004) Chabrier G, Barman T, Baraffe I, Allard F Hauschildt PH (2004) The Evolution of Irradiated Planets: Application to Transits. ApJ 603(1):L53–L56
@@
- Demory and Seager (2011) Demory BO Seager S (2011) Lack of Inflated Radii for Kepler Giant Planet Candidates Receiving Modest Stellar Irradiation. ApJS 197(1):12
- Dos Santos (2021) Dos Santos LA (2021) Observations of planetary winds and outflows. Proc IAU 17(S370):56–71
@@
- Fortney and Nettelmann (2010) Fortney JJ Nettelmann N (2010) The Interior Structure, Composition, and Evolution of Giant Planets. Space Science Reviews 152:423–447
@@
- Fuentes et al. (2022) Fuentes JR, Cumming A Anders EH (2022) Layer formation in a stably stratified fluid cooled from above: Towards an analog for Jupiter and other gas giants. Phys Rev Fluids 7(12):124,501
@@
- Haldemann et al. (2020) Haldemann J, Alibert Y, Mordasini C Benz W (2020) AQUA: A Collection of H$_2$O Equations of State for Planetary Models. A&A 643:A105
@@
- Huang and Cumming (2012) Huang X Cumming A (2012) Ohmic Dissipation in the Interiors of Hot Jupiters. ApJ 757(1):47
@@
- Jackson et al. (2008) Jackson B, Greenberg R Barnes R (2008) Tidal Evolution of Close-in Extrasolar Planets. ApJ 678(2):1396
@@
- Kao et al. (2018) Kao MM, Hallinan G, Pineda JS, Stevenson D Burgasser A (2018) The Strongest Magnetic Fields on the Coolest Brown Dwarfs. ApJS 237(2):25
@@
- Lanza (2008) Lanza AF (2008) Hot Jupiters and stellar magnetic activity. A&A 487(3):1163–1170
@@
- Leconte and Chabrier (2012) Leconte J Chabrier G (2012) A new vision of giant planet interiors: Impact of double diffusive convection. A&A 540:A20
- Leconte et al. (2010) Leconte J, Chabrier G, Baraffe I Levrard B (2010) Is tidal heating sufficient to explain bloated exoplanets? Consistent calculations accounting for finite initial eccentricity. A&A 516:A64
- Leconte et al. (2011) Leconte J, Lai D Chabrier G (2011) Distorted, nonspherical transiting planets: Impact on the transit depth and on the radius determination. A&A 528:A41
@@
- Lee et al. (2022) Lee EJ, Karalis A Thorngren DP (2022) Creating the Radius Gap without Mass Loss. The Astrophysical Journal 941:186
@@
- Love (1909) Love AEH (1909) The yielding of the earth to disturbing forces. Proc R Soc Lond A 82(551):73–88
@@
- Mankovich and Fuller (2021) Mankovich CR Fuller J (2021) A diffuse core in Saturn revealed by ring seismology. Nat Astron 5(11):1103–1109
@@
- Marley et al. (2007) Marley MS, Fortney JJ, Hubickyj O, Bodenheimer P Lissauer JJ (2007) On the Luminosity of Young Jupiters. ApJ 655(1):541
@@
- May et al. (2021) May EM, Komacek TD, Stevenson KB et al. (2021) Spitzer Phase-curve Observations and Circulation Models of the Inflated Ultrahot Jupiter WASP-76b. AJ 162(4):158
@@
- Mugrauer and Michel (2020) Mugrauer M Michel KU (2020) Gaia search for stellar companions of TESS Objects of Interest. Astron Nachr 341(10):996–1030
@@
- Naef et al. (2001) Naef D, Latham DW, Mayor M et al. (2001) HD 80606 b, a planet on an extremely elongated orbit. A&A 375(2):L27–L30
- Narang et al. (2023) Narang M, Oza AV, Hakim K et al. (2023) uGMRT observations of the hot-Saturn WASP-69b: Radio-Loud Exoplanet–Exomoon Survey II (RLEES II). Monthly Notices of the Royal Astronomical Society 522(2):1662–1668
@@
- Parmentier et al. (2016) Parmentier V, Fortney JJ, Showman AP, Morley C Marley MS (2016) TRANSITIONS IN THE CLOUD COMPOSITION OF HOT JUPITERS. ApJ 828(1):22
- Paxton et al. (2013) Paxton B, Cantiello M, Arras P et al. (2013) Modules for Experiments in Stellar Astrophysics (MESA): Planets, Oscillations, Rotation, and Massive Stars. ApJS 208(1):4
- Piso et al. (2015) Piso AMA, Youdin AN Murray-Clay RA (2015) Minimum Core Masses for Giant Planet Formation With Realistic Equations of State and Opacities. ApJ 800(2):82
@@
- Poppenhaeger (2019) Poppenhaeger K (2019) How stars and planets interact: A look through the high-energy window. Astron Nachr 340(4):329–333
@@
- Reiners and Christensen (2010) Reiners A Christensen UR (2010) A magnetic field evolution scenario for brown dwarfs and giant planets. A&A 522:A13
- Ricker et al. (2014) Ricker GR, Winn JN, Vanderspek R et al. (2014) Transiting Exoplanet Survey Satellite. J Astron Telesc Instrum Syst 1(1):014,003
@@
- Sainsbury-Martinez et al. (2019) Sainsbury-Martinez F, Wang P, Fromang S et al. (2019) Idealised simulations of the deep atmosphere of hot Jupiters: Deep, hot adiabats as a robust solution to the radius inflation problem. A&A 632:A114
@@
- Schrijver et al. (1989) Schrijver CJ, Cote J, Zwaan C Saar SH (1989) Relations between the photospheric magnetic field and the emission from the outer atmospheres of cool stars. I - The solar CA II K line core emission. ApJ 337:964
- Serenelli and Basu (2010) Serenelli AM Basu S (2010) Determining the Initial Helium Abundance of the Sun. ApJ 719(1):865
@@
- Shkolnik et al. (2003) Shkolnik E, Walker GAH Bohlender DA (2003) Evidence for Planet-induced Chromospheric Activity on HD 179949. ApJ 597(2):1092–1096
- Shkolnik et al. (2005) Shkolnik E, Walker GAH, Bohlender DA, Gu PG Kurster M (2005) Hot Jupiters and Hot Spots: The Short- and Long-Term Chromospheric Activity on Stars with Giant Planets. ApJ 622(2):1075–1090
- Showman and Polvani (2011) Showman AP Polvani LM (2011) Equatorial Superrotation on Tidally Locked Exoplanets. ApJ 738(1):71
- Sing et al. (2019) Sing DK, Lavvas P, Ballester GE et al. (2019) The Hubble Space Telescope PanCET Program: Exospheric Mg ii and Fe ii in the Near-ultraviolet Transmission Spectrum of WASP-121b Using Jitter Decorrelation. AJ 158(2):91
@@
- Thorngren et al. (2019b) Thorngren DP, Marley MS Fortney JJ (2019b) An Empirical Mass–Radius Relation for Cool Giant Planets. Res Notes AAS 3(9):128
@@
- Valsecchi et al. (2014) Valsecchi F, Rasio FA Steffen JH (2014) From Hot Jupiters to Super-Earths via Roche Lobe Overflow. ApJL 793(1):L3
@@
- Vazan et al. (2018) Vazan A, Ormel CW, Noack L Dominik C (2018) Contribution of the core to the thermal evolution of sub-Neptunes. arXiv:181102588 [astro-ph]
@@
- Wong et al. (2021) Wong I, Shporer A, Zhou G et al. (2021) TOI-2109: An Ultrahot Gas Giant on a 16 hr Orbit. AJ 162(6):256
@@
- Yee et al. (2021) Yee SW, Winn JN Hartman JD (2021) How Complete Are Surveys for Nearby Transiting Hot Jupiters? AJ 162(6):240
@@
- Zarka (1998) Zarka P (1998) Auroral radio emissions at the outer planets: Observations and theories. J Geophys Res 103(E9):20,159–20,194